# EFKAN: A KAN-Integrated Neural Operator For Efficient Magnetotelluric Forward Modeling

### v1

@@ 3.4 Prediction by the EFKAN Trained by the Small-Scale TrData
The performance of neural networks is typically limited by the amount of the training data. However, the larger the training data, the higher the computational cost. In this section, we decrease the number of samples in TrData to 5000 to assess the generalization ability of EFKAN. Fig. 20 displays the training loss on the small-scale TrData and the testing loss on the TeData-I. The training of EFNO is early stopped at the 108-th epoch, while EFKAN is terminated at the 149-th epoch with a smaller loss value. The testing loss curves show that EFKAN achieves a smaller loss than EFNO on TeData-I. As shown in Table 6, EFKAN achieves the lowest average $\ell_{1}$-norm error across all three testing datasets, demonstrating its strong generalization ability.
@@
We randomly sample a resistivity model (Fig. 21) from each of the three datasets TeData-I, TeData-II, and TeData-III to show the effectiveness of EFKAN. Fig. 22 presents the predicted apparent resistivity and phase for the smooth resistivity (Fig. 21(a)), and we can observe that both EFNO and EFKAN obtain solutions with high precision. The $\ell_{1}$-norm error by EFNO and EFKAN are close to each other except for the error of $\rho_{xy}$, where the amplitude of the error by EFNO is much higher than that of EFKAN. The 1-D profiles at 0.049 Hz and 10 Hz (Fig. 23) further demonstrate that EFKAN can fit the ground truth better than EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0240, 0.0278, 0.0298, and 0.0321, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0088, 0.0123, 0.0163, and 0.0185, respectively.
@@
As shown in Fig. 24, both EFNO and EFKAN obtain satisfactory results, though the resistivity with rectangular anomalies is not included in TrData. Additionally, the errors are relatively minor and are mainly concentrated in the low-frequency. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.1396, 0.1588, 0.0933, and 0.1231, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.1161, 0.1425, 0.0702, and 0.1160, respectively. In Fig. 25, we present the 1-D profiles, and we can observe that the predictions by EFKAN meet the ground truth well between -100 and -35 km at 0.049 Hz (Fig. 25(a1) to (a4)), and it exhibits superior performance to EFNO at 10 Hz (Fig. 25(b1) to (b4)).
@@
We evaluate the performance for frequencies that are completely different from TrData. The example is shown in Fig. 21, which also includes rectangular anomalies. As shown in Fig. 26, both EFNO and EFKAN are able to provide reasonable apparent resistivity and phase. The errors corresponding to phase are more pronounced than those of apparent resistivity, as well as their distribution. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0844, 0.0797, 0.0911, and 0.1402, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0529, 0.0767, 0.0683, and 0.1403, respectively. From the 1-D profiles (Fig. 27), it can be observed that both EFNO and EFKAN can essentially fit the apparent resistivity and phase at 0.005 Hz as well as the phase at 12.589 Hz. Furthermore, the apparent resistivity predicted by the two methods is very close to the ground truth, yet EFKAN is more accurate than EFNO at 12.589 Hz (Fig. 27(b1) and (b3)).
@@
Figure 20: (a) The training loss of EFNO and EFKAN on the small-scale TrData dataset; (b) The testing loss of EFNO and EFKAN on the TeData-I dataset.
Table 6: Quantitative Results for the Reduced-Scale TrData
Model TeData-I TeData-II TeData-III
@@
Figure 21: Examples for demonstrating the effectiveness of EFKAN trained on the small-scale TrData: (a) The smooth resistivity model randomly sampled from TeData-I; (b) The smooth resistivity model with rectangular anomalies sampled from TeData-II; (c) The smooth resistivity model with rectangular anomalies sampled from TeData-III.
@@ 3.5 Computational Cost
From Tables 4, 5, and 6, it is evident that EFKAN requires approximately 0.01 seconds to predict the EM solutions for a single resistivity model. In contrast, the traditional and widely-used numerical algorithm FDM takes about 10 seconds per resistivity model for forward modeling. This indicates that the computational speed of EFKAN is nearly 1000 times faster than that of FDM. Additionally, EFKAN consumes slightly more time per resistivity model compared to EFNO, with training times per epoch being 93 seconds and 114 seconds for EFNO and EFKAN, respectively. This discrepancy may arise from the fact that the KAN employs a separate function for each variable in the input, and the standard B-spline functions utilized in the KAN are not optimized for parallel computing on GPUs. To enhance computational efficiency, one could consider replacing the B-spline function with alternative univariate functions, such as radial basis functions (RBFs) (Li, 2024), wavelets (Bozorgasl and Chen, 2024), and Jacobi polynomials (SS et al., 2024).
@@ 4 Discussion
The numerical results indicate that EFKAN achieves higher accuracy in apparent resistivity and phase measurements compared to the EFNO (Peng et al., 2022) that consists of FNO and MLP. We believe that is due to the advantage of KAN for science tasks. Based on the Kolmogorov-Arnold representation theorem, KAN has no linear weights at all–every weight parameter is replaced by a univariate function (Liu et al., 2024), such as B-spline functions. Due to the locality and adjustable number of grids in univariate function, KANs can achieve a certain degree of dynamic network architecture and continual learning (Yu et al., 2024). In reference (Yu et al., 2024), Yu et al. (2024) proved that KAN significantly outperforms MLP in symbolic formula representation, while MLP generally outperforms KAN in machine learning, computer vision, natural language processing, and audio processing. Therefore, KANs have demonstrated promising performance in many science tasks including Poisson equation, knot theory, and Anderson localization.
@@
Furthermore, we believe that KAN can endow EFKAN with a certain degree of interpretability. The Kolmogorov-Arnold representation theorem posits that any multivariate continuous function can be represented as a superposition of continuous functions of a single variable. Thus, it is easy to visualize and understand these learnable activation functions, allowing for a more intuitive grasp of how inputs are transformed into outputs, examples are shown in Fig. 2. This design not only overcomes the limitations of fixed activation functions in terms of interpretability and training efficiency but also aligns well with the global and continuous nature of the problems KAN are designed to solve, such as partial differential equations and operator learning tasks. The parameterization of these functions and their learnable nature make KAN more transparent and understandable, compared to the black-box nature of traditional MLPs.
@@
From Fig. 4, Fig. 12, and Fig. 20, we can observer that the testing loss of both EFNO and ENKAN is lower than the training loss, which is contrary to that in many deep learning tasks. By analyzing the training and testing dataset, we speculate that is due to the to the number of samples in the test dataset being significantly fewer than the number of samples in the training dataset. Specifically, the training dataset TrData contains 15000 samples, which is 150 times the number of samples of test dataset TeData-I, TeData-II and TeData-III. Therefore, the accumulated error on the dataset with more samples may be greater than the accumulated error on the dataset with fewer samples. It is worth mentioning that this phenomenon has also occurred in our previous work (Peng et al., 2022). We also find this phenomenon existing in the Fashion MNIST classification, which consisting of a training set of 60000 examples and a test set of 10000 examples. It can be observed that regardless of whether an MLP or CNN is used, the validation loss (i.e., the testing loss in our experiments) is lower than the training loss (https://nvsyashwanth.github.io/machinelearningmaster/fashion-mnist/).
@@ 5 Conclusion
In this study, we present the development of a novel neural operator, EFKAN, for rapid MT forward modeling by extending the FNO with KAN. Within the EFKAN framework, we employ the FNO as the branch network to map the resistivity model to apparent resistivity and phase in the frequency domain, and we replace the MLPs with the KAN to explore the feasibility which utilizes the KAN to map these apparent resistivity and phase to the desired frequencies and locations. Furthermore, we use the spectral method to obtain the smooth and stochastically varying resistivity model, rather than merely embedding anomalies within a homogeneous half-space underground. We conduct nine numerical experiments to validate the effectiveness of EFKAN. The experimental results demonstrate that EFKAN not only achieves higher accuracy compared to EFNO equipped with MLPs but also exhibits faster computational speed compared to the conventional numerical computational method FDM, which suggests that EFKAN could serve as a potential surrogate model for rapid MT forward modeling.
@@ 6 Acknowledgments
Code availability The source codes are available for downloading at the link: https://github.com/linfengyu77/EFKAN.

### v2

@@ 3.4 Performance on the Downscaling Training Dataset
The performance of neural networks is typically limited by the amount of training data. However, the larger the training data, the higher the computational cost. In this section, we reduce the number of samples in the original training dataset to 5000 to assess the generalizability of EFKAN. Fig. 20 displays the training loss on the downscaling original training dataset and the testing loss on testing dataset A. The training of EFNO is stopped early at the 108-th epoch, while EFKAN is terminated at the 149-th epoch with a lower loss value. The testing loss curves show that EFKAN achieves a smaller loss than EFNO on testing dataset A. As shown in Table 6, EFKAN achieves the lowest average $\ell_{1}$-norm error across the three different testing datasets, demonstrating its strong generalizability.
@@
We randomly select a resistivity model (Fig. 21) from each of testing dataset A, B, and C to show the effectiveness of EFKAN. Fig. 22 presents the predicted apparent resistivity and phase for the smooth resistivity (Fig. 21(a)), and we can observe that both EFNO and EFKAN obtain solutions with high precision. The $\ell_{1}$-norm error by EFNO and EFKAN are close to each other except for the error of $\rho_{xy}$, where the amplitude of the error by EFNO is much higher than that of EFKAN. The 1-D profiles at 0.049 Hz and 10 Hz (Fig. 23) further demonstrate that EFKAN can fit the ground truth better than EFNO. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0240, 0.0278, 0.0298, and 0.0321, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0088, 0.0123, 0.0163, and 0.0185, respectively.
@@
As shown in Fig. 24, both EFNO and EFKAN obtain satisfactory results, though the resistivity with rectangular anomalies is not included in the original training dataset. Additionally, the errors are relatively minor and are mainly concentrated in the low frequency. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.1396, 0.1588, 0.0933, and 0.1231, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.1161, 0.1425, 0.0702, and 0.1160, respectively. In Fig. 25, we present the 1-D profiles, and we can observe that the predictions by EFKAN meet the ground truth well between -100 and -35 km at 0.049 Hz (Fig. 25(a1) to (a4)), and it exhibits superior performance to EFNO at 10 Hz (Fig. 25(b1) to (b4)).
@@
We also evaluate the performance for the frequencies that are different from the original training dataset. An example is shown in Fig. 21, which also includes rectangular anomalies. As shown in Fig. 26, both EFNO and EFKAN are able to provide reasonable apparent resistivity and phase. The errors corresponding to phase are more pronounced than those of apparent resistivity, as well as their distribution. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFNO is 0.0844, 0.0797, 0.0911, and 0.1402, respectively. The relative $\ell_{1}$-norm error of $\rho_{xy},\rho_{yx},\phi_{xy}$, and $\phi_{yx}$ by EFKAN is 0.0529, 0.0767, 0.0683, and 0.1403, respectively. From the 1-D profiles (Fig. 27), it can be observed that both EFNO and EFKAN can essentially fit the apparent resistivity and phase at 0.005 Hz as well as the phase at 12.589 Hz. Furthermore, the apparent resistivity predicted by the two methods is very close to the ground truth, yet EFKAN is more accurate than EFNO at 12.589 Hz (Fig. 27(b1) and (b3)).
@@
Figure 20: (a) The training loss of EFNO and EFKAN on the downscaling original training dataset dataset; (b) The testing loss of EFNO and EFKAN on testing dataset A.
Table 6: Quantitative Results for the downscaling original training dataset
Model Testing dataset A Testing dataset B Testing dataset C
@@
Figure 21: Examples for demonstrating the effectiveness of EFKAN trained on the downscaling original training dataset: (a) The smooth resistivity model randomly sampled from testing dataset A; (b) The smooth resistivity model with rectangular anomalies sampled from testing dataset B; (c) The smooth resistivity model with rectangular anomalies sampled from testing dataset C.
@@ 3.5 Computational Cost
From tables 4, 5, and 6, it is evident that EFKAN requires approximately 0.01 seconds to predict the electromagnetic field for a single resistivity model. In contrast, the traditional and widely-used numerical algorithm FDM takes about 10 seconds per resistivity model for forward modeling. This indicates that the computational speed of EFKAN is nearly 1000 times faster than that of FDM. Additionally, EFKAN consumes slightly more time per resistivity model compared to EFNO, with training time per epoch being 93 seconds and 114 seconds for EFNO and EFKAN, respectively. This discrepancy may arise from the fact that the KAN employs a separate function for each variable in the input, and the standard B-spline functions utilized in the KAN are not optimized for parallel computing on GPUs. To enhance computational efficiency, one could consider replacing the B-spline function with alternative univariate functions, such as radial basis functions (RBFs) (Li, 2024), wavelets (Bozorgasl and Chen, 2024), and Jacobi polynomials (SS et al., 2024).
@@ 4 Discussion
The experimental results prove that EFKAN can achieve higher accuracy in apparent resistivity and phase measurements compared to the EFNO (Peng et al., 2022) that powered by FNO and MLP. We believe that this is due to the advantage of KAN for science. Based on the Kolmogorov-Arnold representation theorem, KAN has no linear weights at all–every weight parameter is replaced by a univariate function (Liu et al., 2024), such as B-spline functions. Due to the locality and the adjustable number of grids in the univariate function, KANs can achieve a certain degree of dynamic network architecture and continuous learning (Yu et al., 2024b). Reference (Yu et al., 2024b) proved that KAN significantly outperforms MLP in symbolic formula representation, while MLP generally outperforms KAN in machine learning, computer vision, natural language processing, and audio processing. Therefore, KANs have demonstrated promising performance in many science tasks including Poisson equation, knot theory, and Anderson localization.
@@
Furthermore, we believe that KAN can endow EFKAN with a certain degree of interpretability. The Kolmogorov-Arnold representation theorem posits that any multivariate continuous function can be represented as a superposition of continuous functions of a single variable. Thus, it is feasible to visualize and understand these learnable activation functions, allowing for a more intuitive grasp of how inputs are transformed into output; examples are shown in Fig. 2. This characterization not only overcomes the limitations of fixed activation functions in terms of interpretability and training efficiency but also aligns well with the global and continuous nature of the problems KAN are designed to solve, such as partial differential equations and operator learning tasks. The parameterization of these functions and their learnable nature make KAN more transparent and understandable, compared to the black-box nature of traditional MLPs.
From Fig. 4, Fig. 12, and Fig. 20, we can observe that EFKAN training converging rapidly and stably, while EFNO training is fluctuating and stopped early. That proves that the combination of KAN and FNO does not introduce new issues in parameter tuning, and it can improve training stability and convergence. The computational intensity of forward modeling significantly constrains the efficiency of MT inversion. The iterative inversion algorithm, such as the nonlinear conjugate gradient (Newman and Alumbaugh, 2000), can obtain the inversion result with high resolution by minimizing the misfit between the observed and simulated electromagnetic field. However, the high computation costs required for MT forward modeling restrict exploration of the model parameter space and limit the number of feasible iterations (Bozdağ et al., ). EFKAN shows two significant advantages for further artificial intelligence-driven MT inversion: (1) By significantly increasing the forward modeling speed to reduce the computational cost of inversion, enabling more extensive searches of the model parameter space within a limited time; (2) By improving the accuracy of forward modeling to optimize search directions within the model parameter space, enhancing the precision of inverted resistivity model. Thus, we believe that EFKAN will improve the efficiency of traditional MT inversion methods, and it can provide a new tool to further enhance the generalization capability of conventional deep learning-based MT inversion, such as the CNN-based MT inversion approach (Yu et al., 2025).
The generalization capability of deep learning is largely based on the quality and amount of training datasets. To obtain high-quality training data, we use GRF to generate the training data to reflect the essential characteristics of real conductivity, though it may not fully capture all the complexities present in real-world resistivity distributions, such as sharp discontinuities or fault-like structures. This is important for real applications because it is very difficult or expensive to obtain a large amount of real data to train neural networks. From the experimental results, we can see that EFKAN can achieve good performance on the unseen data (e.g., the rectangular anomalies and the frequencies not present in the training dataset) and satisfactory predictions. Therefore, we consider that the proposed method can provide flexibility enabling researchers to customize their own training sets and rapidly deploy the neural networks for specific MT forward modeling tasks.
@@
In addition, we can observe that the testing loss of both EFNO and ENKAN is lower than the training loss (Fig. 4, Fig. 12, and Fig. 20), which is contrary to that in many deep learning tasks. By analyzing the training and testing dataset, we speculate that this is due to the fact that the number of samples in the test dataset is significantly fewer than the number of samples in the training dataset. Specifically, the original training dataset contains 15000 samples, which is 150 times the number of samples of testing dataset A, B and C. Therefore, the accumulated error on the dataset with more samples may be greater than the accumulated error on the dataset with fewer samples. It is worth mentioning that this phenomenon has also occurred in our previous work (Peng et al., 2022). We also find this phenomenon existing in the Fashion MNIST classification, which consists of a training set of 60000 samples and a test set of 10000 samples. It can be observed that regardless of whether an MLP or CNN is used, the validation loss (that is, the testing loss in our experiments) is lower than the training loss (https://nvsyashwanth.github.io/machinelearningmaster/fashion-mnist/).
@@ 5 Conclusion
In this study, we present the development of a new neural operator for the rapid forward modeling of MT by combining FNO with KAN. We utilize FNO as the branch network to transform the resistivity model to the apparent resistivity and phase in the frequency domain followed by using KAN to map these apparent resistivity and phase at the desired frequencies and/or locations. In addition, we adopt the spectral method to generate smooth and stochastically varying resistivity models as the training data, rather than embedding anomalies within a homogeneous half-space underground. To evaluate the generalization capability of the proposed method, we implement several challenging experiments, including predicting the high-resolution solution using the model trained on low-resolution data and the solution for the unseen data. Experimental results demonstrate that the proposed method has good generalization capability for these challenging tasks, which can achieve high accuracy and is much faster than the traditional numerical computational algorithm in terms of computational speed, suggesting that EFKAN could serve as a potential surrogate model for rapid MT forward modeling.
@@ 6 Acknowledgments
Code availability The source codes are available for download at the link: https://github.com/linfengyu77/EFKAN.